# Online conformal inference for multi-step time series forecasting

### v1

@@ 5 Experiments
We examine the empirical performance of the proposed multi-step conformal prediction methods using two simulated data settings and two real data examples. Throughout the experiments, we adhere to the following parameter settings:
@@
    we focus on the target coverage level $0.91-\alpha=$ 0.9;
@@
    following Angelopoulos et al. (2023), we use a step size parameter of $\gamma=0.005$ for the MACP method, a Theta model as the scorecaster in the MPID method, and a learning rate of $\eta=0.01\hat{B}_{t}$ for quantile tracking in the MPID and AcMCP methods, where $\hat{B}_{t}=\max\{s_{t-\Delta+1|t-\Delta-h+1},\dots,s_{t|t-h}\}$ is the highest score over a tailing window and the window length $\Delta$ is set to be same as the length of the calibration set;
@@
    we adopt a nonlinear saturation function defined as $r_{t}(x)=K_{1}\tan\left(x\log(t)/(tC_{\text{sat }})\right)$, where $\tan(x)=\operatorname{sign}(x)\cdot\infty$ for $x\notin[-\pi/2,\pi/2]$, and constants $C_{\text{sat }},K_{\mathrm{I}}>0$ are chosen heuristically, as suggested by Angelopoulos et al. (2023);
@@
    we consider a clipped version of MACP by imputing infinite intervals with the largest score seen so far.
@@ 5.1 Simulated linear autoregressive process
We first consider a simulated stationary time series which is generated from a simple AR$(2)$ process
@@
where $\varepsilon_{t}$ is white noise with error variance $\sigma^{2}=1$. After an appropriate burn-in period, we generate $N=5000$ data points. Under the sequential split and online learning settings, we create training sets $\mathcal{D}_{\text{tr}}$ and calibration sets $\mathcal{D}_{\text{cal}}$, each with a length of $500$. We use AR$(2)$ models to generate $1$- to $3$-step-ahead point forecasts (i.e. $H=3$), using the Arima() function from the forecast R package (Hyndman et al. 2024). The goal is to generate prediction intervals using various proposed conformal prediction methods and evaluate whether they can achieve the nominal long-run coverage for each separate forecast horizon.
Figure 2: AR(2) simulation results showing rolling coverage, mean and median interval width for each forecast horizon. The displayed curves are smoothed over a rolling window of size $500$. The black dashed line indicates the target level of $0.91-\alpha=$ 0.9.
Figure 2 presents the rolling coverage and interval width of each method for each forecast horizon, with metrics computed over a rolling window of size $500$. The analytic (and optimal) intervals obtained from the AR model are also shown. We observe that MPID and AcMCP achieve approximately the desired $90\%$ coverage level over the rolling windows, while other methods, including the AR model, undergo much wider swings away from the desired level, showing high coverage volatility over time. The trajectories of the rolling mean and median of the interval widths for each method are largely consistent. AcMCP constructs narrower prediction intervals than MPID, despite both methods achieving similar coverage. Moreover, we see that AcMCP tends to offer adaptive prediction intervals, and results in wider intervals especially when competing methods undercover, which is to be expected. In short, AcMCP intervals offer greater adaptivity and more precise coverage compared to AR, MSCP, MWCP and MACP. However, MPID achieves tight coverage but at the cost of constructing wider prediction intervals. This is due to the inclusion of a second model (scorecaster) which introduces additional variance into the generated prediction intervals. The results can be further elucidated with Figure 3, which presents boxplots of rolling coverage and interval width for each method and each forecast horizon.
@@
The inclusion of the last term $\tilde{e}_{t+h|t}$ in AcMCP should only result in a slight difference compared to the version without this term, which we henceforth refer to as MPI. This is because, the inclusion of $\tilde{e}_{t+h|t}$ aims to capture autocorrelations inherent in multi-step forecast errors and focuses on the mean of forecast errors, whereas the whole update of AcMCP operates on quantiles of scores. To illustrate the subtle difference in their results and explore their origins, we visualize their prediction intervals over a truncated period of length $500$, as shown in Figure 4. We observe that AcMCP and MPI indeed construct similar prediction intervals so their lower and upper bounds mostly overlap with each other. The main differences occur around the time 1320 and during the period 1470-1500, where AcMCP tends to have a fanning-out effect, increasing the interval width as the forecast horizon increases, compared to MPI. Figure 4 also presents the prediction interval bounds given by MACP. The prediction intervals of both AcMCP and MACP can capture certain patterns in the actual observations, and there is no consistent pattern indicating dominance of one method over the other in terms of interval width.
@@ 5.2 Simulated nonlinear autoregressive process
where $x_{1,t}$ and $x_{2,t}$ are uniformly distributed on $[0,1]$, and $\varepsilon_{t}$ is white noise with error variance $\sigma^{2}=0.1$. Thus, the time series $y_{t}$ nonlinearly depends on its lagged values $y_{t-1}$ and $y_{t-2}$, as well as exogenous variables $x_{1,t}$ and $x_{2,t}$.
Figure 5: Nonlinear simulation results showing rolling coverage, mean and median interval width for each forecast horizon. The displayed curves are smoothed over a rolling window of size $100$. The black dashed line indicates the target level of $0.91-\alpha=$ 0.9.
@@
After an appropriate burn-in period, we generate $N=2000$ data points. Under the sequential split and online learning settings, we create training sets $\mathcal{D}_{\text{tr}}$ and calibration sets $\mathcal{D}_{\text{cal}}$, each with a length of $500$. Given the nonlinear structure of the DGP, we use feed-forward neural networks with a single hidden layer and lagged inputs to generate $1$- to $3$-step-ahead point forecasts (i.e. $H=3$), using the nnetar() function from the forecast R package (Hyndman et al. 2024). Note that it is not possible to derive analytic intervals for neural networks, thus we do not include neural network models when presenting the results.
Figure 5 illustrates the rolling coverage and interval width of each method, with calculations based on a rolling window of size $100$. We see that MPID and AcMCP are able to maintain minor fluctuations around the target coverage of $90\%$ across all time indices, contrasting with MSCP, MWCP, and MACP, which struggle to sustain the target coverage and display pronounced fluctuations over time. Moreover, all conformal prediction methods, except for MSCP, construct adaptive prediction intervals. They widen intervals in response to undercoverage and narrow them when overcoverage occurs. Notably, MPID and AcMCP demonstrate greater adaptability, displaying higher variability in interval widths compared to competing methods in order to uphold the desired coverage. AcMCP intervals are evidently narrower than MPID intervals for $2$-step-ahead forecasting but wider for $3$-step-ahead forecasting. AcMCP intervals appear to be more reasonable, as they tend to widen with increasing forecast horizons.
@@ 5.3 Electricity demand data
We apply the conformal prediction methods using data comprising daily electricity demand (GW), daily maximum temperature (degrees Celsius), and holiday information for Victoria, Australia, spanning a three-year period from 2012 to 2014. Temperatures are taken from the Melbourne Regional Office weather station in the capital city of Victoria. The left panel of Figure 7 displays the daily electricity demand during 2012–2014, along with temperatures. The right panel shows a nonlinear relationship between electricity demand and temperature, with demand increasing for low temperatures (due to heating) and increasing for high temperatures (due to cooling). The two clouds of points in the right panel correspond to working days and non-working days.
@@
Our response variable is Demand, and we use two covariates: Temperature, and Workday (an indicator variable for if the day was a working day or not). Following Hyndman & Athanasopoulos (2021), we will fit a dynamic regression model with a piecewise linear function of temperature (containing a knot at $18$ degrees) to generate $1$- to $7$-step-ahead point forecasts (i.e. $H=7$). The error series in the regression is assumed to follow an ARIMA model to contain autocorrelation. We use two years of data as training sets to fit dynamic regression models, and use $100$ data points for calibration sets.
Figure 8: Electricity demand data results showing rolling coverage, mean and median interval width for each forecast horizon. The displayed curves are smoothed over a rolling window of size $100$. The black dashed line indicates the target level of $0.91-\alpha=$ 0.9.
@@
Figure 10: Electricity demand data results showing the forecast interval bounds for MACP, MPI, and AcMCP over the whole test set.
We present the results in Figure 8 and Figure 9, comparing the rolling coverage and interval width of each method. These computations are based on a rolling window of size $100$. The DR method corresponds to the analytic intervals obtained from the dynamic regression model. First, we observe that DR consistently achieves a significantly higher coverage than the $90\%$ target coverage, resulting in much wider intervals than other methods for $h=1,2,3,4$. Second, MSCP, MWCP, and MACP fail to sustain the target coverage and noticeably undercover after September 2014 for all forecast horizons, thus leading to narrower intervals than others. Third, MPI, MPID, and AcMCP offer prediction intervals that are wider than those of other conformal prediction methods, effectively mitigating or avoiding the undercoverage issue observed after September 2014. Additionally, we notice that MPID performs slightly worse than MPI and AcMCP in terms of coverage for $h=3,4,7$, despite leading to wider intervals. Finally, MPI and AcMCP coverage display similar pattern, but AcMCP is capable of constructing narrower intervals than MPI.
We present the forecast interval bounds for MACP, MPI, and AcMCP in Figure 10. The plot shows that MACP intervals are too narrow to adequately hug the true values, particularly from September to November 2014. In contrast, MPI and AcMCP perform better by widening their intervals, resulting in narrower swings away from the $90\%$ target level. The differences between MPI and AcMCP prediction intervals are most pronounced in the $5$-, $6$-, and $7$-step-ahead forecast results. For $5$-step-ahead forecasting, from May to July 2014, the upper bounds of MPI intervals struggle to capture the true values. AcMCP addresses this undercoverage by construct larger upper bounds. In August and September, AcMCP constructs smaller upper bounds than MPI while still capturing the true values effectively. For $6$-step-ahead forecasting from May to July, AcMCP offers smaller upper bounds than MPI, which provides upper bounds that are far away from the truth. Similar reaction is observed for $7$-step-ahead forecasting during August and September.
@@ 5.4 Eating out expenditure data
In our final example, we apply the conformal prediction methods to forecast the eating out expenditure (\$million) in Victoria, Australia. The data set includes monthly expenditure on cafes, restaurants and takeaway food services in Victoria from April 1982 up to December 2019, as shown in Figure 11. The data shows an overall upward trend, obvious annual seasonal patterns, and variability proportional to the data level.
@@
We consider three models: ARIMA with logarithmic transformation, ETS, and STL-ETS (Hyndman & Athanasopoulos 2021), and then output their simple average as final point forecasts. STL-ETS involves forecasting using the STL decomposition method, applying ETS to forecast the seasonally adjusted series. All three models can be automatically trained using the forecast R package (Hyndman et al. 2024). Our goal is to forecast $12$ months ahead, i.e. $H=12$. We use $20$ years of data for training the models and $5$ years of data for calibration sets. The whole test set only has a length of $152$ months. Therefore, we will not compute rolling coverage and interval width in this experiment, but rather compute the coverage and interval width averaged over the entire test set.
We summarise the average coverage and interval width for each method and each forecast horizon in Figure 12. The results first show that MSCP, MWCP and MACP provide valid prediction intervals for smaller forecast horizon but fail to achieve the desired coverage for larger forecast horizons ($h>5$). Second, for $h\leq 5$, MPI and AcMCP can approximately achieve the desired coverage and provide comparable mean interval widths with other methods, except for MPID. Third, the coverage of MSCP, MWCP and MACP declines gradually as the forecast horizon increases, while MPI and AcMCP maintain coverage within a tighter range, albeit at the cost of interval efficiency. Lastly, compared to MPI, AcMCP exhibits slightly less deviation from the desired coverage across most forecast horizons.

### v2

@@ 5 Experiments
We evaluate the proposed multi-step conformal prediction methods using two simulated settings and two real-world data sets. The following parameter choices are used throughout:
@@
    we focus on the target coverage level $1-\alpha=0.9$;
@@
    following Angelopoulos et al. (2023), MACP uses a step size $\gamma=0.005$; MPID uses a Theta model as the scorecaster; and both MPID and AcMCP use a learning rate of $\eta=0.01\hat{B}_{t}$ for quantile tracking, where $\hat{B}_{t}=\max\{s_{t-\Delta+1|t-\Delta-h+1},\dots,s_{t|t-h}\}$ is the highest score over a tailing window of length $\Delta$, which is set equal to the calibration set size;
@@
    we adopt a nonlinear saturation function $r_{t}(x)=K_{1}\tan\left(x\log(t)/(tC_{\text{sat }})\right)$, where $\tan(x)=\operatorname{sign}(x)\cdot\infty$ for $x\notin[-\pi/2,\pi/2]$, and constants $C_{\text{sat }},K_{\mathrm{I}}>0$ are chosen heuristically as in Angelopoulos et al. (2023);
@@
    we use a clipped version of MACP, replacing infinite intervals with the largest score observed to date. Other methods do not produce infinite intervals by construction and therefore require no clipping.
@@ 5.1 Simulated linear autoregressive process
We first consider a simulated stationary time series generated from a simple AR(2)(2) process
@@
where $\varepsilon_{t}$ is white noise with error variance $\sigma^{2}=1$. After an appropriate burn-in period, we generate $N=5000$ data points. Under the sequential split and online learning settings, we create training sets $\mathcal{D}_{\text{tr}}$ and calibration sets $\mathcal{D}_{\text{cal}}$, each with a length of $500$. We use AR(2)(2) models to generate $1$- to $3$-step-ahead point forecasts (i.e. $H=3$), using the Arima() function from the forecast R package (Hyndman & Khandakar 2008, Hyndman et al. 2026). The goal is to generate prediction intervals using various proposed conformal prediction methods and evaluate whether they can achieve the nominal long-run coverage for each separate forecast horizon.
Figure 2: AR(2) simulation results showing rolling coverage, mean and median interval width for each forecast horizon. The displayed curves are smoothed over a rolling window of size $500$. The target coverage level is $1-\alpha=0.9$.
Figure 2 presents the rolling coverage and interval width of each method for each forecast horizon, with metrics computed over a rolling window of size $500$. The analytic (and optimal) intervals obtained from the AR model are also shown. We observe that MPID and AcMCP achieve approximately the desired $90\%$ coverage level over the rolling windows, while other methods, including the AR model, undergo much wider swings away from the desired level, showing high coverage volatility over time. For each forecast horizon and method, the rolling mean and median interval widths follow very similar trajectories across rolling windows. AcMCP constructs narrower prediction intervals than MPID, despite both methods achieving similar coverage. Moreover, we see that AcMCP tends to offer coverage-adaptive prediction intervals, automatically adjusting their width based on past coverage performance, and results in wider intervals especially when competing methods undercover, which is to be expected. In short, AcMCP intervals offer greater adaptivity and more precise coverage compared to AR, MSCP, MWCP and MACP. Both MPID and AcMCP achieve tight coverage, but AcMCP outputs more informative and smaller intervals. The benefits of AcMCP are more noticeable when $h$ grows, and it does not deteriorate for smaller $h$. This improvement can be attributed to the inclusion of a second model (scorecaster) which introduces additional variance into the generated prediction intervals. The results can be further elucidated with Figure 3, which presents boxplots of rolling coverage and interval width for each method and each forecast horizon.
@@
The inclusion of the last term $\tilde{e}_{t+h|t}$ in AcMCP should only result in a slight difference compared to the version without this term, which we henceforth refer to as MPI. This is because, the inclusion of $\tilde{e}_{t+h|t}$ aims to capture autocorrelations inherent in multi-step forecast errors and focuses on the mean of forecast errors, whereas the whole update of AcMCP operates on quantiles of scores. To illustrate the subtle difference in their results and explore their origins, we visualize their prediction intervals over a truncated period of length $500$, as shown in Figure 4. We observe that AcMCP and MPI indeed construct similar prediction intervals so their lower and upper bounds mostly overlap with each other. The main differences occur around the time 1320 and during the period 1470-1500, where AcMCP tends to have a fanning-out effect, increasing the interval width as the forecast horizon increases, compared to MPI. Figure 4 also presents the prediction interval bounds given by MACP. Since both MACP and AcMCP rely on past miscoverage information to update prediction intervals, we include MACP to support a direct comparison of their resulting prediction intervals while preserving a clear and concise presentation. The prediction intervals of both AcMCP and MACP can capture certain patterns in the actual observations, and there is no consistent pattern indicating dominance of one method over the other in terms of interval width.
@@ 5.2 Simulated nonlinear autoregressive process
where $x_{1,t}$ and $x_{2,t}$ are uniformly distributed on $[0,1]$, and $\varepsilon_{t}$ is white noise with variance $\sigma^{2}=0.1$. Thus, $y_{t}$ nonlinearly depends on its lagged values $y_{t-1}$, $y_{t-2}$, and exogenous variables $x_{1,t}$ and $x_{2,t}$.
Figure 5: Nonlinear simulation results showing rolling coverage, mean and median interval width for each forecast horizon. The displayed curves are smoothed over a rolling window of size $100$. The target coverage level is $1-\alpha=0.9$.
@@
After an appropriate burn-in period, we generate $N=2000$ data points. Under the sequential split and online learning settings, both training set $\mathcal{D}_{\text{tr}}$ and calibration set $\mathcal{D}_{\text{cal}}$ have size $500$. Given the nonlinear DGP, we use feed-forward neural networks with a single hidden layer and lagged inputs to generate $1$- to $3$-step-ahead point forecasts ($H=3$), using the nnetar() function from the forecast R package (Hyndman et al. 2026). Note that analytic prediction intervals are unavailable for neural networks, thus we do not include them when presenting the results.
Figure 5 shows the rolling coverage and interval width of each method, computed on a rolling window of size $100$. MPID and AcMCP are able to maintain minor fluctuations around the target coverage of $90\%$ over time, whereas MSCP, MWCP, and MACP struggle to sustain the target level and display pronounced fluctuations over time. Moreover, all methods, except for MSCP, adapt interval widths to distributional changes, widening intervals in response to undercoverage and narrowing them under overcoverage. Notably, MPID and AcMCP demonstrate greater adaptability, with higher variability in interval widths compared to competing methods in order to uphold the desired coverage. The outcome of the rolling mean and median interval widths indicates that AcMCP produces generally narrower intervals than MPID for $2$-step-ahead forecasts but wider intervals for $3$-step-ahead forecasts. Furthermore, under AcMCP, the interval width at $h=3$ exceeds that at $h=2$ in $53.41\%$ of the test cases, compared with only $38.08\%$ under MPID. This pattern suggests more plausible behavior for AcMCP, as forecast uncertainty typically increases with the forecast horizon, leading to a greater tendency for wider prediction intervals.
@@ 5.3 Electricity demand data
We apply the conformal prediction methods using data comprising daily electricity demand (GW), daily maximum temperature (degrees Celsius), and holiday information for Victoria, Australia, from 2012 to 2014. Temperatures are taken from the Melbourne Regional Office weather station. The left panel of Figure 7 displays the daily electricity demand and temperatures over the period. The right panel shows a nonlinear relationship between demand and temperature, with demand increasing for low temperatures (due to heating) and increasing for high temperatures (due to cooling); the two clouds of points correspond to working days and non-working days.
@@
The response variable is Demand, with two covariates: Temperature, and Workday (an indicator variable for working days). Following Hyndman & Athanasopoulos (2021), we fit a dynamic regression model with a piecewise linear function of temperature (with a knot at $18$ degrees) to generate $1$- to $7$-step-ahead point forecasts ($H=7$), modeling the errors with an ARIMA process to capture autocorrelation. We use two years of data for training and $100$ observations for calibration.
Figure 8 and Figure 9 compare the rolling coverage and interval width across method. These computations are based on a rolling window of size $100$. The DR method corresponds to the analytic intervals obtained from the dynamic regression model. First, DR consistently attains coverage well above the $90\%$ target, resulting in much wider intervals than other methods for $h=1,2,3,4$. While this overcoverage improves reliability, the resulting intervals are overly conservative and may lead to inefficient operational decisions in electricity markets, such as excessive reserve procurement. Second, MSCP, MWCP, and MACP fail to sustain the target coverage and noticeably undercover after September 2014 for all horizons, thus leading to narrower intervals than others. In electricity market applications, such undercoverage underestimates demand uncertainty and increases the risk of insufficient reserve allocation. Third, MPI, MPID, and AcMCP offer wider intervals that largely mitigate the post–September 2014 undercoverage. Among these, MPID shows slightly poorer coverage than MPI and AcMCP at $h=3$, despite producing to wider intervals. Finally, while MPI and AcMCP display similar coverage pattern, AcMCP is capable of constructing narrower intervals than MPI, particularly at larger horizons. Relative to MPID, AcMCP also achieves greater interval width reduction, with advantages that become more pronounced as $h$ increases and without sacrificing performance at shorter horizons.
Figure 8: Electricity demand data results showing rolling coverage, mean and median interval width for each forecast horizon. The displayed curves are smoothed over a rolling window of size $100$. The target coverage level is $1-\alpha=0.9$.
@@ 5.4 Eating out expenditure data
In our final example, we apply the conformal prediction methods to forecast the eating out expenditure (\$ million) in Victoria, Australia. The data includes monthly expenditure on cafes, restaurants and takeaway food services from April 1982 to December 2019, as shown in Figure 10. The data shows an overall upward trend, obvious annual seasonal, and evel-dependent variability.
@@
We consider three models: ARIMA with logarithmic transformation, ETS, and STL-ETS (Hyndman & Athanasopoulos 2021), and then output their simple average as final point forecasts. STL-ETS applies ETS to the seasonally adjusted series obtained via STL decomposition. All models are automatically trained using the forecast R package (Hyndman et al. 2026). We aim to forecast $12$ months ahead ($H=12$) using $20$ years of data for training and $5$ year for calibration. Because the test set contains only $152$ months, we report coverage and interval width averaged over the full test period rather than rolling windows.
As the forecast horizons considered in this application are relatively long, we summarize performance in Figure 11 by reporting the average coverage gap and average interval width across the whole test set for each method and each forecast horizon. The results first show that MSCP, MWCP and MACP provide valid prediction intervals for smaller forecast horizon but fail to achieve the desired coverage for larger forecast horizons ($h>5$). Second, for $h\leq 5$, MPI and AcMCP can approximately achieve the desired coverage and provide comparable mean interval widths with other methods, except for MPID. Third, the coverage of MSCP, MWCP and MACP declines gradually as the forecast horizon increases, while MPI and AcMCP maintain coverage within a tighter range, albeit at the cost of interval efficiency. Lastly, compared to MPI, AcMCP exhibits slightly less deviation from the desired coverage across most forecast horizons.
@@
5.5 Post-hoc comparison
For MACP, the step size $\gamma$ controls both the calibration tightness and the responsiveness of the interval widths: smaller values lead to smoother interval evolution but slower correction of local coverage errors, while larger values improve responsiveness at the cost of increased variability. An analogous trade-off arises in AcMCP, where the learning rate $\eta$ governs the balance between adaptivity and stability. This naturally raises the question of whether AcMCP’s empirical advantages persist when the benchmark MACP is optimally tuned for high adaptivity.
To examine this, we conduct a post-hoc comparison in which $\eta$ for AcMCP is held fixed, while $\gamma$ for MACP is selected via grid search over $[0.005,0.5]$. Specifically, $\gamma$ is chosen to match the sample variance of local coverage achieved by AcMCP over rolling windows of the test set.
Table 1 summarizes, for each forecast horizon, the mean coverage over the full test set, the minimum and maximum local coverage across rolling windows, and the mean and median interval widths for MACP and AcMCP across the data sets considered above. In simple linear and nonlinear simulations, both methods attain mean coverage close to the nominal 90%, with AcMCP consistently producing comparable or narrower intervals when coverage is comparable. More pronounced differences arise in real data sets and at longer horizons. For electricity demand, MACP displays systematic undercoverage as the horizon increases, particularly in terms of minimum local coverage, whereas AcMCP maintains more stable coverage around the target by adaptively adjusting interval widths. A similar pattern appears for eating out expenditure data, where MACP suffers substantial coverage degradation at larger horizons despite tuning, while AcMCP preserves near-nominal coverage, albeit with wider but informative intervals. Overall, these results indicate that AcMCP provides superior coverage control and, when coverage is comparable, yields prediction intervals that are shorter or at least comparable in length.
Table 1: Performance of MACP and AcMCP across the four datasets, summarized by mean test-set coverage, minimum and maximum local coverage over rolling windows, and mean and median interval widths for each forecast horizon.
Mean coverage (%) Min. local coverage (%) Max. local coverage (%) Mean interval width Median interval width
MACP AcMCP MACP AcMCP MACP AcMCP MACP AcMCP MACP AcMCP
AR(2) simulation data
h=1 89.95 90.00 89.20 89.40 90.80 90.80 4.04 3.55 3.62 3.54
h=2 89.89 89.94 89.00 89.00 91.00 91.00 5.08 4.68 4.85 4.66
h=3 89.94 90.02 88.80 88.80 91.00 91.20 5.16 4.68 5.00 4.84
Nonlinear simulation data
h=1 90.00 90.00 87.00 86.00 94.00 93.00 0.35 0.36 0.33 0.36
h=2 89.98 90.18 86.00 87.00 94.00 93.00 0.37 0.37 0.33 0.36
h=3 89.86 90.06 87.00 87.00 94.00 94.00 0.35 0.38 0.33 0.38
Electricity demand data
h=1 89.92 89.15 86.00 85.00 91.00 92.00 25.02 24.44 22.52 24.24
h=2 88.67 90.23 83.00 86.00 91.00 94.00 30.56 30.72 29.03 28.87
h=3 88.58 90.94 81.00 86.00 92.00 93.00 34.81 34.35 31.64 33.57
h=4 89.29 90.87 84.00 88.00 93.00 94.00 41.93 40.74 34.90 39.48
h=5 89.60 94.80 86.00 91.00 93.00 99.00 41.17 55.62 34.72 45.43
h=6 90.32 90.73 85.00 88.00 92.00 97.00 36.34 46.65 35.08 41.56
h=7 87.80 89.84 84.00 87.00 90.00 95.00 36.55 47.45 35.02 47.98
Eating out expenditure data
h=1 85.62 88.89 80.00 83.33 90.00 93.33 76.70 65.51 64.21 63.08
h=2 84.77 90.01 78.33 83.33 93.33 96.67 97.93 91.50 84.26 89.10
h=3 80.54 89.26 68.33 83.33 95.00 95.00 104.10 112.42 101.46 107.83
h=4 59.18 89.12 15.00 81.67 91.67 98.33 89.74 116.80 77.68 111.16
h=5 46.90 91.03 13.33 81.67 65.00 96.67 59.72 126.12 68.74 120.97
h=6 36.36 90.91 8.33 85.00 41.67 96.67 29.26 178.17 62.05 178.60
h=7 53.90 86.52 31.67 73.33 61.67 98.33 63.10 186.58 81.44 198.27
h=8 47.48 88.29 23.33 83.33 58.33 98.33 55.30 192.07 78.62 192.59
h=9 32.12 89.78 1.67 86.67 48.33 100.00 20.29 194.40 47.47 168.96
h=10 71.85 89.63 53.33 85.00 90.00 93.34 116.88 195.83 121.56 193.68
h=11 69.92 88.72 41.67 85.00 98.33 96.67 102.20 233.80 99.26 220.92
h=12 69.47 91.60 65.00 83.33 83.33 98.33 127.19 336.02 128.28 297.48